Find the lateral area and volume of the frustum of a right circular cone having a slant height of 14m and the radii of the bases are 4m and 1.5 m respectively.


1
Expert's answer
2021-12-21T15:44:33-0500

"A=\pi(r1+r2)\times l\\ HERE \,
r1=4,r2=1.5,l=14\\=\pi(4+1.5)\times 14\\\pi(5.5)\times 14\\=241.90m^2"


volume

"=\frac{\pi H}{3}(\frac{r1^3-r2^3}{3})\\=295.08m^3"